# Experimental evidence of production of directional muons from a laser-wakefield accelerator

### v3

@@ I. INTRODUCTION
High-power lasers have been recently proposed as a viable alternative to produce sizeable muon fluxes in a relatively compact setup (see, e.g. Refs. [9, 10, 11, 12]), with numerical work showing the possibility of achieving high-energy populations of muons with an average flux sufficient for practical applications with a turnaround as short as a few seconds to minutes [10]. Muons can be generated during the interaction of GeV-scale laser-wakefield accelerated electron beams with thick high-Z converters through several channels: Bethe-Heitler pair production [13], trident processes [14], pion decay [15], and kaon decay [16]. At GeV energies, the high mass of kaons makes its production comparatively unlikely. While pion-mediated muon production has the highest cross-section [17], the wide-angle emission of charged pions, combined with the large cross sections of pion scattering and absorption in heavy nuclei, results in generally not collimated muon fluxes [15]. For sufficiently thick targets, Bethe-Heitler pair production dominates over direct electroproduction (trident) processes [14, 15], making it the preferable mechanism for generating sizeable, directional muon fluxes from a laser-wakefield accelerator [10, 11, 15]. Bethe-Heitler pair production has been used to generate laser-driven ultra-relativistic positron beams (see, e.g., Refs. [18, 19, 20, 21, 22]). However, the much higher rest mass of muons results in significantly lower yields compared to electron-positron pairs (). This results in a high degree of complexity in muon detection, due to the co-presence of a high level of noise arising from electron-positron pairs and bremsstrahlung photons. Several schemes have been numerically proposed to physically separate laser-driven muon beams from other secondary particles (see, e.g., Ref. [10]).
Here, we report on experimental evidence of the production of directional muons driven by a PW-scale laser system. Statistical analysis of the experimental data indicates detection of muon signal with a % confidence over noise, in close agreement with numerical simulations. Numerical extrapolation of these results to a PW-scale 10 Hz laser system such as EPAC in the UK [23] indicates that up to muons/s can be isolated and guided onto cm2-scale areas. This configuration will allow, for instance, for the high resolution detection of strategically sensitive materials in high-Z thick containers within minutes [10].
@@ III. NUMERICAL MODELLING
A 2 cm-thick Pb converter was used, since this is the thickness expected to maximise the flux of high-energy directional muons [10]. Positive and negative muons present a similar spectrum (Fig. 2.a) extending approximately up to 1.25 GeV and an energy-dependent divergence of the order of a hundred mrad (consistently with simulation results reported in Ref. [10]). Simulations indicate a total of and per primary electron. For an initial electron beam charge of pC, this translates into and escaping the converter target per shot.
@@
A major obstacle in the detection of muons is the high-flux of scattered secondary particles exiting the converter. Numerical simulations indicate the generation of a total of scattered electrons, bremsstrahlung photons, and generated positrons, with spectra extending up to the maximum energy of the primary electron beam (see also [19, 20]) and charged pions with a broad angular distribution. In order to minimise the signal induced by these secondary particles in the detector, a specific shielding configuration has been designed and implemented in the experiment (Fig. 1.a). In this configuration, it is preferable to measure the positive muons, where the noise is lower. Monte-Carlo FLUKA simulations of the entire experimental configuration have been carried out to extract the number and spectral properties of muons and noise particles reaching the detector plane (1.41.4 cm2, as dictated by the size of the Timepix3 detector). The simulation results (see Fig. 3) indicate approximately muons, electrons, positrons, gammas, and pions per primary electron reaching the detector per shot. The muons generated via pion decay represent a negligible fraction of the muons reaching the detector (%, see Figs. 2.b and 3.b). Due to the magnetic transport line, the muons reaching the detector plane have an energy of () MeV; the other noise particles (i.e., electrons, positrons, and gammas) have instead a broad energy spectrum up to approximately 10 MeV.
@@
The muon signal can be statistically distinguished from other particles by the characteristics of energy deposition in the detector, a standard technique in particle physics experiments (see, e.g., Ref. [33]). Although methods based on measuring muon lifetime are also possible in principle (see, e.g., Ref. [12]), the high electromagnetic noise generated in the chamber by the high-power laser in this experiment prevented a clear signal identification using this technique. High-energy muons produce distinctive track shapes and energy deposition patterns on the detector, distinguishing them from MeV-scale leptonic noise (electrons hereafter for definitiveness), ions, and neutrons. While pions could also in principle contribute to the signal [34], the simulation results discussed above indicate that only 0.070.02 pions reach the detector per shot. This low likelihood arises from their wide-angle emission combined with the magnetic transport line and the lead shielding in front of the detector (see Fig. 3.f). Our simulations also confirmed that Timepix3 based on thin silicon detectors are also inefficient in detecting MeV-scale photons, with negligible energy deposition (see also Ref. [35]).
@@
Ions and neutrons are expected to arrive at the detector plane significantly after the temporal gating applied to the detector (discussed below) and can thus be easily discarded by time-filtering the signal. For muons and electrons, the distributions of the mean and standard deviation of the energy deposited per track and the length of the tracks are shown in Fig. 4. In all cases, Kolmogorov-Smirnov tests of the distributions indicated a -value < 0.001 confirming that the distributions are statistically different. Muon tracks are generally shorter, with a higher mean and standard deviation of the energy deposited. This is accompanied by a different shape of the tracks; muon tracks have negligible curvature and are never observed to be more than 1 pixel wide. This is to be intuitively expected, since high-energy muons will have much lower scattering angles in the detector, thus resulting in narrower and straighter energy deposition patterns.
@@ IV. EXPERIMENTAL RESULTS AND ANALYSIS
The filtered data still show approximately 30 tracks, which could in principle be produced by either a muon or an electron. In order to further distinguish between electron and muon signal, likelihood ratios (LRs) for the length of the tracks and the mean and standard deviation of the energy deposited are shown in Fig. 6, with brown dots representing each track in Fig. 5.d. The corresponding histogram of the combined LR for each track in Fig. 5.d is shown in orange in Fig. 7. An LR above 3 (below 0.3) is considered substantial evidence of muon (electron) behaviour, while an LR above 10 (below 0.1) is considered strong evidence [36]. The histogram is well approximated by a two-Gaussian distribution with means at and , providing strong preliminary indication that most tracks are produced by electrons. Notably, three tracks (labelled A,B, and C in Figs. 6 and 7) have combined LRs above 3 ( and ), first preliminary indication of being produced by a muon.
For a more quantitative analysis, the LRs have been used to extract the probability that each track in Fig. 5 is produced by a muon or an electron. For this, the pre-test probability has been calculated by applying the same filters of time of arrival, straightness, and track width to the simulation data. After this process, simulations indicate electron tracks and muon tracks over ten shots, yielding a pre-test probability %. and pre-test odds %. The post-test odds have thus been calculated for each track using the combined likelihood ratio: . The probability that at least one muon has been detected can thus be calculated as:
@@ V. CONCLUSIONS
The results reported here can be readily scaled to laser systems of high peak power or repetition rate. Extrapolation to a nC-scale 2 GeV electron beam, as attainable by PW-scale lasers [37], indicates that up to muons per shot can be isolated and guided by a suitable transport line [10]. A 10 Hz PW laser such as EPAC in the UK [23] would then be able to produce and guide up to muons/s in shielded cm2-scale areas, enabling radiography of extended high-Z objects within tens of seconds [10]. 10 PW-scale laser facilities are also now starting to enter the international landscape [38] (see, e.g., Vulcan 2020 [39], OPAL [40], and ELI-NP [24]). Numerical simulations show that the 10PW line at ELI-NP can produce multi-nC electron beams with multi-GeV maximum energies [10]. The interaction of such electron beams with a 2 cm converter is predicted to produce more than muons with a broadband energy distribution peaking at 0.7 GeV. A suitable transport line can thus conservatively direct up to muons at the sample plane in a 10x10 cm2 area [10]. Even operating at one shot per minute, 10 PW-scale lasers would enable muon radiography of extended high-Z materials to be carried out within a few minutes [10].
Acknowledgments: G.S. wishes to acknowledge support from EPSRC (grant No. EP/V044397/1, EP/V049186/1, and 2934194) and EU Horizon Europe R&I Grant No. 101079773 (EuPRAXIA-PP). P.M. wishes to acknowledge support from EPSRC (grant No. EP/V049232/1). B.B. and P.S. have profited from funding from the Czech Science Foundation under Registration Number GM23-04869M. R.E.M. has been supported by the Global Postdoc Fellowship Program of the Czech Technical University in Prague. The authors wish to acknowledge support from the Vulcan dark period community support programme 23-1 from the Rutherford Appleton Laboratory. This work was supported by the Extreme Light Infrastructure–Nuclear Physics (ELI-NP) Phase II, a project co-financed by the Romanian Government and the European Union through the European Regional Development Fund, by the Romanian Ministry of Education and Research CNCS-UEFISCDI (Project No. PN-IIIP4-IDPCCF-2016-0164) and Nucleu Projects Grant No. PN 23210105. The Romanian government also supports ELI-NP through IOSIN funds as a Facility of National Interest.

### v4

@@ I. INTRODUCTION
High-power lasers have been recently proposed as a viable alternative to produce sizeable muon fluxes in a relatively compact setup (see, e.g. Refs. [28, 7]), with numerical work showing the possibility of achieving high-energy populations of muons with an average flux sufficient for practical applications with a turnaround as short as a few seconds to minutes [7]. Proof-of-principle experimental evidence of muon generation driven by PW-scale lasers has been independently reported [35, 40], where muon signal was identified by their signature decay time.
Muons can be generated during the interaction of GeV-scale laser-wakefield accelerated electron beams with thick high-Z converters through several channels: Bethe-Heitler pair production [38], trident processes [3], pion decay [37], and kaon decay [29]. At GeV energies, the high mass of kaons makes its production comparatively unlikely. While pion-mediated muon production has the highest cross-section [32], the wide-angle emission of charged pions, combined with the large cross sections of pion scattering and absorption in heavy nuclei, results in generally not collimated muon fluxes [37]. For sufficiently thick targets, Bethe-Heitler pair production dominates over direct electroproduction (trident) processes [3, 37], making it the preferable mechanism for generating sizeable, directional muon fluxes from a laser-wakefield accelerator [7, 35, 37]. Bethe-Heitler pair production has been used to generate laser-driven ultra-relativistic positron beams (see, e.g., Refs. [34, 31, 30, 8, 21]). However, the much higher rest mass of muons results in significantly lower yields compared to electron-positron pairs (). This results in a high degree of complexity in muon detection, due to the co-presence of a high level of noise arising from electron-positron pairs and bremsstrahlung photons. Several schemes have been numerically proposed to physically separate laser-driven muon beams from other secondary particles (see, e.g., Ref. [7]).
Here, we report on experimental evidence of the production of high-energy and directional muons. For the first time, the experimental setup allows selecting high-energy muons and provides a pathway for the separation of muons from leptonic and photon noise, validating a novel detection mechanism and recent numerical modelling. The detected muons are generated by the Bethe-Heitler process in a converter material irradiated by laser-wakefield accelerated GeV-scale electron beams. These muons exit the converter in a directional manner, with a divergence of the order of a 100 mrad. Statistical analysis of the experimental data indicates detection of muon signal with a % confidence over noise, in close agreement with numerical simulations. Numerical extrapolation of these results to a PW-scale 10 Hz laser system such as EPAC in the UK [16] indicates that up to muons/s can be isolated and guided onto cm2-scale areas. This configuration will allow, for instance, for the high resolution detection of strategically sensitive materials in high-Z thick containers within minutes [7].
@@ III. NUMERICAL MODELLING
Extensive numerical modelling of muon production in a similar experimental configuration [7] indicates that the yield of high-energy directional muons increases linearly with target thickness up to approximately 2 cm. For thicker targets, the muon yield only marginally increases with the additional drawback of an enlarged divergence and, thus, decrease in flux. While optimum thickness will ultimately depend on the specific electron beam parameters and the muon population of interest, this thickness is optimum for this experimental configuration and is thus used hereafter.
@@
Positive and negative muons present a similar spectrum (Fig. 2.a) extending approximately up to 1.25 GeV and an energy-dependent divergence of the order of a hundred mrad (consistently with simulation results reported in Ref. [7]). Given the nature of the muon generation process, the muons exiting the converter target are expected to present a broadband, bremsstrahlung-like, spectrum virtually irrespective of the spectrum of the primary electron beam.
Simulations indicate a total of and per primary electron. For an initial electron beam charge of pC, this translates into and escaping the converter target per shot.
@@
A major obstacle in the detection of muons is the high-flux of scattered secondary particles exiting the converter. Numerical simulations indicate the generation of a total of scattered electrons, bremsstrahlung photons, and generated positrons, with spectra extending up to the maximum energy of the primary electron beam (see also [31, 30]) and charged pions with a broad angular distribution. In order to minimise the signal induced by these secondary particles in the detector, a specific shielding configuration has been designed and implemented in the experiment (Fig. 1.a). In this configuration, it is preferable to measure the positive muons, where the noise is lower. Monte-Carlo FLUKA simulations of the entire experimental configuration have been carried out to extract the number and spectral properties of muons and noise particles reaching the detector plane (1.41.4 cm2, as dictated by the size of the Timepix3 detector). The simulation results (see Fig. 3) indicate approximately muons, electrons, positrons, gammas, and pions per primary electron reaching the detector per shot. The muons generated via pion decay represent a negligible fraction of the muons reaching the detector (%, see Figs. 2.b and 3.b). Due to the magnetic transport line, the muons reaching the detector plane have an energy of () MeV; the other noise particles reaching the detector (i.e., electrons, positrons, and gammas) have instead a broad energy spectrum up to approximately 10 MeV (Fig 4).
@@
The muon signal can be statistically distinguished from other particles by the characteristics of energy deposition in the detector, a standard technique in particle physics experiments (see, e.g., Ref. [1]). Although methods based on measuring muon lifetime are also possible [35, 40], the high electromagnetic noise generated in the chamber by the high-power laser in this experiment prevented a clear signal identification using this technique. High-energy muons produce distinctive track shapes and energy deposition patterns on the detector, distinguishing them from MeV-scale leptonic noise (electrons hereafter for definitiveness), ions, and neutrons. While pions could also in principle contribute to the signal [33], the simulation results discussed above indicate that only 0.070.02 pions reach the detector per shot. This low likelihood arises from their wide-angle emission combined with the magnetic transport line and the lead shielding in front of the detector (see Fig. 3.f). Our simulations also confirmed that Timepix3 based on thin silicon detectors are also inefficient in detecting MeV-scale photons, with negligible energy deposition (see also Ref. [18]).
@@
Ions and neutrons are expected to arrive at the detector plane significantly after the temporal gating applied to the detector (discussed below) and can thus be easily discarded by time-filtering the signal. For muons and electrons, the distributions of the mean and standard deviation of the energy deposited per track and the length of the tracks are shown in Fig. 5. In all cases, Kolmogorov-Smirnov tests of the distributions indicated a -value < 0.001 confirming that the distributions are statistically different. Muon tracks are generally shorter, with a higher mean and standard deviation of the energy deposited. This is accompanied by a different shape of the tracks; muon tracks have negligible curvature and are never observed to be more than 1 pixel wide. These numerical results are consistent with the significance difference in energy deposition characteristics between high-energy muons and low energy electrons or positrons, with the latter experiencing larger scattering angles and a more continuous energy transfer throughout the material.
@@ IV. EXPERIMENTAL RESULTS AND ANALYSIS
The filtered data still show approximately 30 tracks, which could in principle be produced by either a muon or an electron. In order to further distinguish between electron and muon signal, likelihood ratios (LRs) for the length of the tracks and the mean and standard deviation of the energy deposited are shown in Fig. 7, with brown dots representing each track in Fig. 6.d. The likelihood ratio that a track has a property (for example, length) given that it is generated by a particle type and not any other particle () is defined as , where the probabilities are extracted from the distributions shown in Fig. 5. A track can be parameterised by its length , the mean energy deposited , and its standard deviation . The combined likelihood for a given track is then the product . The corresponding histogram of the combined LR for each track in Fig. 6.d is shown in orange in Fig. 8. An LR above 3 (below 0.3) is considered substantial evidence of muon (electron) behaviour, while an LR above 10 (below 0.1) is considered strong evidence [19]. The histogram is well approximated by a two-Gaussian distribution with means at and , providing strong preliminary indication that most tracks are produced by electrons. Notably, three tracks (labelled A,B, and C in Figs. 7 and 8) have combined LRs above 3 ( and ), first preliminary indication of being produced by a muon.
For a more quantitative analysis, the LRs have been used to extract the probability that each track in Fig. 6 is produced by a muon or an electron. For this, the pre-test probability has been calculated by applying the same filters of time of arrival, straightness, and track width to the simulation data. After this process, simulations indicate electron tracks and muon tracks over ten shots, yielding a pre-test probability %, and pre-test odds for a track to be a muon. The equivalent pre-test probability and odds for a track to not be a muon, i.e. is an electron, are and respectively. The post-test odds that a given track, , is of particle type can thus be calculated as: . The probability that each track has been generated by a muon can thus be expressed as , with the probability of it being generated by an electron being necessarily . Given the fact that each track can be treated independently, the probability that at least one muon has been detected can thus be calculated as:
@@ V. CONCLUSIONS
The results reported here can be readily scaled to laser systems of high peak power or repetition rate. Extrapolation to a nC-scale 2 GeV electron beam, as attainable by PW-scale lasers [25], indicates that up to muons per shot can be isolated and guided by a suitable transport line [7]. A 10 Hz PW laser such as EPAC in the UK [16] would then be able to produce and guide up to muons/s in shielded cm2-scale areas, enabling radiography of extended high-Z objects within tens of seconds [7]. 10 PW-scale laser facilities are also now starting to enter the international landscape [9] (see, e.g., Vulcan 2020 [17], OPAL [14], and ELI-NP [12]). Numerical simulations show that the 10PW line at ELI-NP can produce multi-nC electron beams with multi-GeV maximum energies [7]. The interaction of such electron beams with a 2 cm converter is predicted to produce more than muons with a broadband energy distribution peaking at 0.7 GeV. The increase in muon yield expected with PW-class laser systems can also be accompanied by a more complex transport line and shielding configuration as proposed and numerically studied in [7]. Adopting an experimental configuration of that kind can thus conservatively direct up to muons in a relatively background-free region with a 10x10 cm2 area [7]. Even operating at one shot per minute, 10 PW-scale lasers would enable muon radiography of extended high-Z materials to be carried out within a few minutes [7].
Acknowledgments: G.S. wishes to acknowledge support from EPSRC (grant No. EP/V044397/1, EP/V049186/1, and 2934194) and EU Horizon Europe R&I Grant No. 101079773 (EuPRAXIA-PP). P.M. wishes to acknowledge support from EPSRC (grant No. EP/V049232/1). B.B. and P.S. acknowledge funding from the Czech Science Foundation under Registration Number GM23-04869M. R.E.M. has been supported by the Global Postdoc Fellowship Program of the Czech Technical University in Prague. The authors wish to acknowledge support from the Vulcan dark period community support programme 23-1 from the Rutherford Appleton Laboratory. This work was supported by the Extreme Light Infrastructure–Nuclear Physics (ELI-NP) Phase II, a project co-financed by the Romanian Government and the European Union through the European Regional Development Fund, by the Romanian Ministry of Education and Research CNCS-UEFISCDI (Project No. PN-IIIP4-IDPCCF-2016-0164) and Nucleu Projects Grant No. PN 23210105. The Romanian government also supports ELI-NP through IOSIN funds as a Facility of National Interest.